The length of a rectangle is 5ft longer than its width.
If the perimeter of the rectangle is 30ft, find its area.

Answer by checkley77(12844)

L=W+5
2L+2W=30
2(W+5)+2W=30
2W+10+2W=30
4W=30-10
4W=20
W=20/4
W=5 FT. ANS. FOR THE WIDTH
L=5+5=10 FT. ANS. FOR THE LENGTH.
AREA=LW
AREA=10*5=50 FT^2 ANS.